# LancBiO: Dynamic Lanczos-aided Bilevel Optimization via Krylov Subspace

### v1

@@ 1 Introduction
Bilevel optimization, in which upper-level and lower-level problems are nested with each other, mirrors a multitude of applications, e.g., game theory (Stackelberg, 1952), hyper-parameter optimization (Franceschi et al., 2017, 2018), meta-learning (Bertinetto et al., 2018), neural architecture search (Liu et al., 2018; Wang et al., 2022), adversarial training (Wang et al., 2021), reinforcement learning (Chakraborty et al., 2023; Hong et al., 2023).
In this paper, we consider the following bilevel problem:
@@
where the upper-level function $f$ and the lower-level function $g$ are defined on $\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}$. $\varphi$ is called the hyper-objective, and the gradient of $\varphi(x)$ is referred to as the hyper-gradient (Pedregosa, 2016; Grazzi et al., 2020; Chen et al., 2023; Yang et al., 2023) if it exists. In contrast to standard single-level optimization problems, bilevel optimization is inherently challenging due to its intertwined structure. Specifically, the formulation (1) underscores the crucial role of the lower-level solution $y^{*}(x)$ in each update of $x$.
One of the focal points in recent bilevel methods has shifted towards nonconvex upper-level problems coupled with strongly convex lower-level problems (Ghadimi & Wang, 2018; Ji et al., 2021; Chen et al., 2022; Dagréou et al., 2022; Li et al., 2022; Hong et al., 2023). This configuration ensures that $y^{*}(x)$ is a single-valued function of $x$, i.e., $y^{*}(x)=\operatorname*{arg\,min}_{y\in\mathbb{R}^{d_{y}}}g(x,y)$. Subsequently, the hyper-gradient can be computed from the implicit function theorem as follows (Ghadimi & Wang, 2018),
$\nabla\varphi(x)=$ $~{}\nabla_{x}f\left(x,y^{*}\left(x\right)\right)-\nabla^{2}_{xy}g\left(x,y^{*}\left(x\right)\right)$ (2)
×$\times\left[\nabla_{yy}^{2}g\left(x,y^{*}\left(x\right)\right)\right]^{-1}\nabla_{y}f\left(x,y^{*}\left(x\right)\right).$
The gradient methods based on the hyper-gradient, $x_{k+1}=x_{k}-\lambda\nabla\varphi(x_{k}),$ are known as the approximate implicit differentiation (AID) based methods (Ji et al., 2021; Arbel & Mairal, 2022; Liu et al., 2023). Nevertheless, the computation of the hyper-gradient (2) suffers from two pains: 1) solving the lower-level problem to obtain $y^{*}(x)$; 2) assembling the Hessian inverse vector product
@@
To this end, it is beneficial to adopt a few inner iterations to approximate $y^{*}(x)$ and $v^{*}(x)$ within each outer iteration (i.e., the update of $x$). Note that the approximation accuracy of $v^{*}(x)$ is crucial for the AID-based methods; see (Ji et al., 2022; Li et al., 2022). Specifically, Figure 1 confirms that the more inner iterations, the higher quality of the estimate of $v^{*}$, and the more enhanced descent of the objective function within the same number of outer iterations..
Figure 1: Test loss for the method stocBiO (Ji et al., 2021) with different inner iterations $I$ to approximate the Hessian inverse vector product.
Approximation: Existing efforts are dedicated to approximating $v^{*}$ in different fashions by regulating the number of inner iterations, e.g., the Neumann series approximation (Ghadimi & Wang, 2018; Ji et al., 2021) for the inverse, gradient descent (Arbel & Mairal, 2022; Dagréou et al., 2022) and conjugate gradient descent (Pedregosa, 2016; Yang et al., 2023) for the linear system.
Amortization: Moreover, there are studies aimed at amortizing the cost of approximation through outer iterations. These methods include using the inner estimate from the previous outer iteration as a warm start for the current outer iteration (Ji et al., 2021; Arbel & Mairal, 2022; Dagréou et al., 2022; Ji et al., 2022; Li et al., 2022; Xiao et al., 2023), or employing a refined step size control (Hong et al., 2023).
@@ 1.1 Contributions
Moreover, the constructed subspaces enable us to dynamically and incrementally approximate $v^{*}$ across outer iterations, thereby achieving an enhanced estimate of the hyper-gradient; Figure 2 illustrates that the proposed LancBiO reaches the best estimation error for $v^{*}$. Hence, we provide a new perspective for approximating the Hessian inverse vector product in bilevel optimization. Specifically, the number of Hessian-vector products averages at $(1+\frac{1}{m})$ per outer iteration with the subspace dimension $m$, which is favorably comparable with the existing methods.
Figure 2: Estimation error of the Hessian inverse vector product in hyper-data cleaning task with corruption rate $0.5$ for different methods: LancBiO and SubBiO (ours), AmIGO (Arbel & Mairal, 2022), and SOBA (Dagréou et al., 2022).
@@ 1.2 Related Work
Krylov subspace methods: Subspace techniques have gained significant recognition in the realm of numerical linear algebra (Parlett, 1998; Saad, 2011; Golub & Van Loan, 2013) and nonlinear optimization (Yuan, 2014; Liu et al., 2021). Specifically, numerous optimization methods utilized subspace techniques to improve efficiency, including acceleration technique (Li et al., 2020), diagonal preconditioning (Gao et al., 2023), and derivative-free optimization methods (Cartis & Roberts, 2023). Krylov subspace (Krylov, 1931), due to its special structure,
@@
with the dimension $N$ for a matrix $A$ and a vector $b$, exhibits advantageous properties in convex quadratic optimization (Nesterov et al., 2018), eigenvalue computation (Kuczyński & Woźniakowski, 1992), and regularized nonconvex quadratic problems (Carmon & Duchi, 2018). Krylov subspace has been widely considered in large-scale optimization such as trust region methods (Gould et al., 1999), trace maximization problems (Liu et al., 2013), and cubic Newton methods (Cartis et al., 2011; Jiang et al., 2024). Lanczos process (Lanczos, 1950) is an orthogonal projection method onto the Krylov subspace, which reduces a dense symmetric matrix to a tridiagonal form. Details of the Krylov subspace and the Lanczos process are summarized in Appendix B.
Approximation of the Hessian inverse vector product : It is cumbersome to compute the Hessian inverse vector product in bilevel optimization. To bypass it, several strategies implemented through inner iterations were proposed, e.g., the Neumann series approximation (Ghadimi & Wang, 2018; Ji et al., 2021), gradient descent (Arbel & Mairal, 2022; Dagréou et al., 2022), and conjugate gradient descent (Pedregosa, 2016; Arbel & Mairal, 2022; Yang et al., 2023). Alternatively, the previous information was exploited in (Ji et al., 2021; Arbel & Mairal, 2022) as a warm start for outer iterations; Ramzi et al. (2022) suggested approximating the Hessian inverse in the manner of quasi-Newton; Dagréou et al. (2022) and Li et al. (2022) proposed the frameworks without inner iterations to approximate the Hessian inverse vector product.
@@ 2 Subspace-based Algorithms
The (hyper-)gradient descent method carries out the $k$-th outer iteration as follows,
$x_{k+1}=x_{k}-\lambda\nabla\varphi(x_{k}),$
where the hyper-gradient is exactly computed by
$\nabla\varphi(x_{k})=\nabla_{x}f\left(x_{k},y^{*}_{k}\right)-\nabla^{2}_{xy}g\left(x_{k},y^{*}_{k}\right)v^{*}_{k}$
with $y^{*}_{k}:=y^{*}(x_{k})$ and $v^{*}_{k}:=v^{*}(x_{k})$ defined in (3). In view of the computational intricacy of $y^{*}_{k}$ and $v^{*}_{k}$, it is commonly concerned with the following estimator for the hyper-gradient
@@
where $y_{k}$ is an approximation of $y^{*}_{k}$. Denote the Hessian
$A_{k}=\nabla_{yy}^{2}g(x_{k},y_{k})\quad\text{and}\quad b_{k}=\nabla_{y}f(x_{k},y_{k}).$
$v_{k}$ is the (approximate) solution of a quadratic optimization problem
@@
Can we construct a low-dimensional subspace $\mathcal{S}_{k}$ such that the solution of (6) satisfactorily approximates $A_{k}^{-1}b_{k}$?
General subspace constructions introduced in the existing subspace methods (Yuan, 2014; Liu et al., 2021) are not straightforward and not exploited in the bilevel setting, rendering the exploration of appropriate subspaces challenging. In the following subsections, we construct approximate Krylov subspaces and propose an elaborate subspace-based framework for bilevel problems.
@@ 2.1 Why Krylov subspace: the SubBiO algorithm
In light of the Neumann series for a suitable $\eta\in\mathbb{R}$
$A^{-1}b=\eta\sum_{i=0}^{\infty}(I-\eta A)^{i}b,$ (7)
it is observed from Appendix B that $A^{-1}b$ belongs to a Krylov subspace for some $N>0$, i.e.,
@@
Given a constant $n\ll N$, we consider an approximation of $A^{-1}b$ in a lower-dimensional Krylov subspace $\mathcal{K}_{n}(A,b)$, i.e., $v_{n}\in\mathcal{K}_{n}(A,b)=\mathcal{K}_{n}(I-\eta A,b)$ and
$v_{n}=\sum_{i=0}^{n-1}{c_{i}\left(I-\eta A\right)^{i}b}\approx A^{-1}b.$
Note that the approximation $v_{n}$ is composed of the set $\{\left(I-\eta A\right)^{i}b\}_{i=0}^{n-1}$ in the sense of the Neumann series (7). Moreover, we observe that $(I-\eta A)v_{n}\in\mathcal{K}_{n+1}(A,b)$ and hence we can recursively choose
@@ 2.2 Why dynamic Lanczos: the LancBiO framework
where $S_{k}:=[b_{k}\,\,\left(I-\eta A_{k}\right)v_{k-1}]\in\mathbb{R}^{d_{y}\times 2}$. The solution $z^{*}\in\mathbb{R}^{2}$ results in $v_{k}=S_{k}z^{*}$. It is a two-dimensional subproblem, whereas computing the projection $S_{k}^{\top}A_{k}S_{k}$ requires two Hessian-vector products, which dominate the cost of the subproblem. Therefore, it is crucial to reduce or amortize the projection cost while preserving the advantages of the Krylov subspace. To this end, we find that the Lanczos process (Appendix B) provides an enlightening way (Lanczos, 1950; Saad, 2011; Golub & Van Loan, 2013) since it allows for the construction of a Krylov subspace and maintaining a tridiagonal matrix as the projection matrix, which significantly reduces the computational cost.
In bilevel optimization, since the quadratic problem (6) evolves through outer iterations, it is difficult to leverage the Lanczos process to amortize the projection cost while simultaneously updating variables as in SubBiO. Specifically, the Lanczos process is inherently unstable (Paige, 1980), and thus the accumulative difference among $\{A_{k}\}$ and $\{b_{k}\}$ will make the Lanczos process invalid.
@@
Consequently, $Q_{k}$ approximates the basis of the true Krylov subspace $\mathcal{K}_{k}(A_{k},b_{k})$, and $T_{k}$ approximates the projection of $A_{k}$ onto it, i.e.,
$T_{k}\approx Q_{k}^{\top}A_{k}Q_{k}.$
However, $Q_{k}$ will lose the orthogonality due to the evolution of $A_{j}$, and $T_{k}$ will deviate from the true projection. Based on this observation, we restart the subspace spanned by the matrix $Q$ for every $m$ outer iterations (line 3 to line 10 in Algorithm 2). The restart mechanism allows us to mitigate the accumulation of the difference among $\{A_{1},\dots,A_{k}\}$, and hence, we can maintain a more reliable basis matrix to approximate Krylov subspaces. The above dynamic Lanczos subroutine, DLanczos, is summarized in Appendix C (line 11 in Algorithm 2).
Residual minimization: The preceding discussion reveals that the dimension of the subspace $\mathcal{S}_{k}$ should be moderate to retain the reliability of the dynamic process. Nevertheless, the limited dimension stagnates the approximation of the subproblem (6) to the full space problem. Therefore, instead of solving the quadratic subproblem (6) in the subspace, we intend to find $v\in\mathcal{S}_{k}=\operatorname{span}(Q_{k})$ such that $A_{k}v\approx b_{k}$. To this end, after going through $m$ outer iterations, we denote the current approximation by $\bar{v}$. In the subsequent outer iterations, we concentrate on a linear system with a residual in the form
@@
and use the solution $\Delta v_{k}$ as a correction to $\bar{v}$ (line 12 to line 14 in Algorithm 2),
$v_{k}=\bar{v}+\Delta v_{k}.$
@@
The combination of the two strategies, restart mechanism and residual minimization, not only controls the dimension of the subspace but also utilizes historical information to enhance the approximation accuracy. By considering a simplified scenario, we reduce the two strategies into solving a standard linear system problem $Ax=b$ with $A$ and $b$ fixed. Note that, from the perspective of Theorem 1 in (Carmon & Duchi, 2018), the residual associated with solving a linear system in an $m$-dimensional Krylov subspace decays faster than a rate $\mathcal{O}\left({{1}/{m^{2}}}\right)$ after each restart. In other words, the estimation error of the Hessian inverse vector product experiences a decay rate of $\mathcal{O}\left({{1}/{m^{2}}}\right)$ after every restart, i.e.,
$\frac{\left\|b-Av_{m\left(h+1\right)}\right\|}{\left\|b-Av_{mh}\right\|}=\frac{\left\|\left(b-A\bar{v}_{h}\right)-A\Delta_{m\left(h+1\right)}\right\|}{\left\|b-A\bar{v}_{h}\right\|}=\mathcal{O}\left(\frac{1}{m^{2}}\right).$
@@
The classic Lanczos process is known for its capability to solve indefinite linear systems (Greenbaum et al., 1999). In the same fashion, the LancBiO framework can be adapted to the bilevel problems with a nonconvex lower-level problem. Interested readers are referred to Appendix D for more details.
2.3 Relation to the existing algorithms
The proposed SubBiO and LancBiO have intrinsic connections to the existing algorithms. Generally, in each outer iteration, methods such as BSA (Ghadimi & Wang, 2018) and TTSA (Hong et al., 2023) truncate the Neumann series (7) at $N$, exploiting information from an $N$-dimensional Krylov subspace. On the contrary, both SubBiO and LancBiO implicitly gather knowledge from a high-dimensional Krylov subspace with less effort.

### v2

@@ 1 Introduction
Bilevel optimization, in which upper-level and lower-level problems are nested with each other, mirrors a multitude of applications, e.g., game theory (Stackelberg, 1952), hyper-parameter selection (Ye et al., 2023), data poisoning (Liu et al., 2024), meta-learning (Bertinetto et al., 2018), neural architecture search (Liu et al., 2018; Wang et al., 2022), adversarial training (Wang et al., 2021), reinforcement learning (Chakraborty et al., 2024; Thoma et al., 2024; Yang et al., 2025), computer vision (Liu et al., 2021a). In this paper, we consider the bilevel problem:
@@
where the upper-level function $f$ and the lower-level function $g$ are defined on $\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}$. $\varphi$ is called the hyper-objective, and the gradient of $\varphi(x)$ is referred to as the hyper-gradient (Pedregosa, 2016; Grazzi et al., 2020; Yang et al., 2023; Chen et al., 2024) if it exists. In contrast to standard single-level optimization problems, bilevel optimization is inherently challenging due to its intertwined structure. Specifically, the formulation (1) underscores the crucial role of the lower-level solution $y^{*}(x)$ in each update of $x$.
One of the focal points in recent bilevel methods has shifted towards nonconvex upper-level problems coupled with strongly convex lower-level problems (Ghadimi and Wang, 2018; Ji et al., 2021; Dagréou et al., 2022; Li et al., 2022; Hong et al., 2023; Hu et al., 2024). This configuration ensures that $y^{*}(x)$ is a single-valued function of $x$, i.e., $y^{*}(x)=\operatorname*{arg\,min}_{y\in\mathbb{R}^{d_{y}}}g(x,y)$. Subsequently, $\nabla\varphi(x)$ can be computed via the implicit function theorem following (Ghadimi and Wang, 2018),
$\nabla\varphi(x)=$ $~{}\nabla_{x}f\left(x,y^{*}\left(x\right)\right)-\nabla^{2}_{xy}g\left(x,y^{*}\left(x\right)\right)\left[\nabla_{yy}^{2}g\left(x,y^{*}\left(x\right)\right)\right]^{-1}\nabla_{y}f\left(x,y^{*}\left(x\right)\right).$ (2)
Gradient methods based on the hyper-gradient, $x_{k+1}=x_{k}-\lambda\nabla\varphi(x_{k}),$ are known as the approximate implicit differentiation (AID) based methods (Ji et al., 2021; Liu et al., 2023b; Huang, 2024). However, the computation of the hyper-gradient (2) suffers from two pains: 1) solving the lower-level problem to obtain $y^{*}(x)$; 2) assembling the Hessian inverse vector product
@@
To this end, it is beneficial to adopt a few inner iterations to approximate $y^{*}(x)$ and $v^{*}(x)$ within each outer iteration (i.e., the update of $x$). Note that the approximation accuracy of $v^{*}(x)$ is crucial for AID-based methods; see (Ji et al., 2022; Li et al., 2022). Specifically, the left of Figure 1 confirms that the more inner iterations, the higher quality of the estimate of $v^{*}$, and the more enhanced descent of the objective function within the same number of outer iterations.
Figure 1: Left: test loss for the method stocBiO (Ji et al., 2021) with different inner iterations $I$ to approximate the Hessian inverse vector product; Right: Estimation error of the Hessian inverse vector product in hyper-data cleaning task with corruption rate $0.5$ for different methods: LancBiO and SubBiO (ours), AmIGO (Arbel and Mairal, 2022), and SOBA (Dagréou et al., 2022).
Approximation: Existing efforts are dedicated to approximating $v^{*}$ in different fashions by regulating the number of inner iterations, e.g., the Neumann series approximation (Ghadimi and Wang, 2018; Ji et al., 2021) for the inverse, gradient descent (Arbel and Mairal, 2022; Dagréou et al., 2022) and conjugate gradient descent (Pedregosa, 2016; Yang et al., 2023) for the linear system.
Amortization: Moreover, there are studies aimed at amortizing the cost of approximation through outer iterations. These methods include using the inner estimate from the previous outer iteration as a warm start for the current outer iteration (Ji et al., 2021; Arbel and Mairal, 2022; Dagréou et al., 2022; Ji et al., 2022; Li et al., 2022; Xiao et al., 2023), or employing a refined step size control (Hong et al., 2023).
@@ 1.1 Contributions
Moreover, the constructed subspaces enable us to dynamically and incrementally approximate $v^{*}$ across outer iterations, thereby achieving an enhanced estimate of the hyper-gradient; the right of Figure 1 illustrates that the proposed LancBiO reaches the best estimation error for $v^{*}$. Hence, we provide a new perspective for approximating the Hessian inverse vector product in bilevel optimization. Specifically, the number of Hessian-vector products averages at $(1+{1}/{m})$ per outer iteration with the subspace dimension $m$, which is favorably comparable with the existing methods.
@@ 1.2 Related Work
Krylov subspace methods: Subspace techniques have gained significant recognition in the realm of numerical linear algebra (Parlett, 1998; Saad, 2011; Golub and Van Loan, 2013) and nonlinear optimization (Yuan, 2014; Liu et al., 2021c). Specifically, numerous optimization methods utilized subspace techniques to improve efficiency, including acceleration technique (Li et al., 2020), diagonal preconditioning (Gao et al., 2023), and derivative-free optimization methods (Cartis and Roberts, 2023). Krylov subspace (Krylov, 1931), due to its special structure,
@@
with the dimension $N$ for a matrix $A$ and a vector $b$, exhibits advantageous properties in convex quadratic optimization (Nesterov et al., 2018), eigenvalue computation (Kuczyński and Woźniakowski, 1992), and regularized nonconvex quadratic problems (Carmon and Duchi, 2018). Krylov subspace has been widely considered in large-scale optimization such as trust region methods (Gould et al., 1999), trace maximization problems (Liu et al., 2013), and cubic Newton methods (Cartis et al., 2011; Jiang et al., 2024). Lanczos process (Lanczos, 1950) is an orthogonal projection method onto the Krylov subspace, which reduces a dense symmetric matrix to a tridiagonal form. Details of the Krylov subspace and the Lanczos process are summarized in Appendix B.
Approximation of the Hessian inverse vector product : It is cumbersome to compute the Hessian inverse vector product in bilevel optimization. To bypass it, several strategies implemented through inner iterations were proposed, e.g., the Neumann series approximation (Ghadimi and Wang, 2018; Ji et al., 2021), gradient descent (Arbel and Mairal, 2022; Dagréou et al., 2022), and conjugate gradient descent (Pedregosa, 2016; Arbel and Mairal, 2022; Yang et al., 2023). Alternatively, the previous information was exploited in (Ji et al., 2021; Arbel and Mairal, 2022) as a warm start for outer iterations; Ramzi et al. (2022) suggested approximating the Hessian inverse in the manner of quasi-Newton; Dagréou et al. (2022) and Li et al. (2022) proposed the frameworks without inner iterations to approximate the Hessian inverse vector product.
@@ 2 Subspace-based Algorithms
The (hyper-)gradient descent method carries out the $k$-th outer iteration as $x_{k+1}=x_{k}-\lambda\nabla\varphi(x_{k})$, where the hyper-gradient is exactly computed by $\nabla\varphi(x_{k})=\nabla_{x}f\left(x_{k},y^{*}_{k}\right)-\nabla^{2}_{xy}g\left(x_{k},y^{*}_{k}\right)v^{*}_{k}$ with $y^{*}_{k}:=y^{*}(x_{k})$ and $v^{*}_{k}:=v^{*}(x_{k})$ defined in (3). In view of the computational intricacy of $y^{*}_{k}$ and $v^{*}_{k}$, it is commonly concerned with the following estimator for the hyper-gradient
@@
where $y_{k}$ is an approximation of $y^{*}_{k}$. Denote $A_{k}=\nabla_{yy}^{2}g(x_{k},y_{k})$ and $b_{k}=\nabla_{y}f(x_{k},y_{k})$. The vector $v_{k}$ serves as the (approximate) solution of the quadratic optimization problem
@@
Can we construct a low-dimensional subspace $\mathcal{S}_{k}$ such that
the solution of the quadratic problem (6) satisfactorily approximates $A_{k}^{-1}b_{k}$?
General subspace constructions introduced in the existing subspace methods (Yuan, 2014; Liu et al., 2021c) are not straightforward and not exploited in the bilevel setting, rendering the exploration of appropriate subspaces challenging. In the following subsections, we construct approximate Krylov subspaces and propose an elaborate subspace-based framework for bilevel problems.
@@ 2.1 Why Krylov subspace: the SubBiO algorithm
In light of the Neumann series for a suitable $\eta\in\mathbb{R}$, $A^{-1}b=\eta\sum_{i=0}^{\infty}(I-\eta A)^{i}b$, it is observed from Appendix B that $A^{-1}b$ belongs to a Krylov subspace for some $N>0$, i.e.,
@@
Given a constant $n\ll N$, we consider an approximation of $A^{-1}b$ in a lower-dimensional Krylov subspace $\mathcal{K}_{n}(A,b)$, i.e., $v_{n}\in\mathcal{K}_{n}(A,b)=\mathcal{K}_{n}(I-\eta A,b)$ and $v_{n}=\sum_{i=0}^{n-1}{c_{i}\left(I-\eta A\right)^{i}b}\approx A^{-1}b$. Note that the approximation $v_{n}$ is composed of the set $\{\left(I-\eta A\right)^{i}b\}_{i=0}^{n-1}$ in the sense of the Neumann series. Moreover, we observe that $(I-\eta A)v_{n}\in\mathcal{K}_{n+1}(A,b)$ and hence we can recursively choose
@@ 2.2 Why dynamic Lanczos: the LancBiO framework
where $S_{k}:=[b_{k}\,\,\left(I-\eta A_{k}\right)v_{k-1}]\in\mathbb{R}^{d_{y}\times 2}$. The solution $z^{*}\in\mathbb{R}^{2}$ results in $v_{k}=S_{k}z^{*}$. It is a two-dimensional subproblem, whereas computing the projection $S_{k}^{\top}A_{k}S_{k}$ requires two Hessian-vector products, which dominate the cost of the subproblem. Therefore, it is crucial to reduce or amortize the projection cost while preserving the advantages of the Krylov subspace. To this end, we find that the Lanczos process (Appendix B) provides an enlightening way (Lanczos, 1950; Saad, 2011; Golub and Van Loan, 2013) since it allows for the construction of a Krylov subspace and maintaining a tridiagonal matrix as the projection matrix, which significantly reduces the computational cost.
In bilevel optimization, since the quadratic problem (6) evolves through outer iterations, it is difficult to leverage the Lanczos process to amortize the projection cost while updating variables as in SubBiO. Specifically, the Lanczos process is inherently unstable (Paige, 1980), and thus the accumulative difference among $\{A_{k}\}$ and $\{b_{k}\}$ will make the Lanczos process invalid.
@@
Consequently, $Q_{k}$ approximates the basis of the true Krylov subspace $\mathcal{K}_{k}(A_{k},b_{k})$, and $T_{k}$ approximates the projection of $A_{k}$ onto it, i.e., $T_{k}\approx Q_{k}^{\top}A_{k}Q_{k}$. However, $Q_{k}$ will lose the orthogonality due to the evolution of $A_{j}$, and $T_{k}$ will deviate from the true projection. Based on this observation, we restart the subspace spanned by the matrix $Q$ for every $m$ outer iterations (line 3 to line 10 in Algorithm 2). The restart mechanism allows us to mitigate the accumulation of the difference among $\{A_{1},\dots,A_{k}\}$, and hence, we can maintain a more reliable basis matrix to approximate Krylov subspaces. The above dynamic Lanczos subroutine, DLanczos, is summarized in Appendix C.
Residual minimization: The preceding discussion reveals that the dimension of the subspace $\mathcal{S}_{k}$ should be moderate to retain the reliability of the dynamic process. However, the limited dimension stagnates the approximation of the subproblem (6) to the full space problem. Therefore, instead of directly solving the quadratic subproblem (6) in the subspace, we intend to find $v\in\mathcal{S}_{k}=\operatorname{span}(Q_{k})$ such that $A_{k}v\approx b_{k}$. To this end, after going through $m$ outer iterations, we denote the current approximation by $\bar{v}$. In the subsequent outer iterations, we concentrate on a linear system with a residual in the form
@@
and use the solution $\Delta v_{k}$ as a correction to $\bar{v}$ (line 12 to line 14 in Algorithm 2), $v_{k}=\bar{v}+\Delta v_{k}.$
@@
The combination of the two strategies, restart mechanism and residual minimization, not only controls the dimension of the subspace but also utilizes historical information to enhance the approximation accuracy. By considering a simplified scenario, we reduce the two strategies into solving a standard linear system problem $Ax=b$ with $A$ and $b$ fixed. Note that, from the perspective of Theorem 1 in (Carmon and Duchi, 2018), the residual associated with solving a linear system in an $m$-dimensional Krylov subspace decays faster than a rate $\mathcal{O}\left({{1}/{m^{2}}}\right)$ after each restart. In other words, the estimation error of the Hessian inverse vector product experiences a decay rate of $\mathcal{O}\left({{1}/{m^{2}}}\right)$ after every restart, i.e.,
$\frac{\left\|b-Av_{m\left(h+1\right)}\right\|^{2}}{\left\|b-Av_{mh}\right\|^{2}}$ $=\frac{\left\|\left(b-A\bar{v}_{h}\right)-A\Delta_{m\left(h+1\right)}\right\|^{2}}{\left\|b-A\bar{v}_{h}\right\|^{2}}$
$=\mathcal{O}\left(\frac{1}{m^{2}}\right).$
@@
The classic Lanczos process is known for its capability to solve indefinite linear systems (Greenbaum et al., 1999). In a similar fashion, the LancBiO framework can be adapted to the bilevel problems with a nonconvex lower-level problem. Interested readers are referred to Appendix D for details.
2.3 Relation to existing algorithms
The proposed SubBiO and LancBiO have intrinsic connections to the existing algorithms. Generally, in each outer iteration, methods such as BSA (Ghadimi and Wang, 2018) and TTSA (Hong et al., 2023) truncate the Neumann series at $N$, exploiting information from an $N$-dimensional Krylov subspace. In contrast, both SubBiO and LancBiO implicitly gather knowledge from a higher-dimensional Krylov subspace with less effort.